OpenStudy (anonymous):

how do I do partial fractions for: (2+s^2)/((s^2)(s+1)(s+2)) ? Thanks for the help!

OpenStudy (anonymous):

As+b c d ----- + --------- + -------- s^2 s+1 s+2

OpenStudy (anonymous):

how do I go from there though?

OpenStudy (anonymous):

set it equal to (2+s^2)/((s^2)(s+1)(s+2))

OpenStudy (anonymous):

then multiply by denominator

OpenStudy (anonymous):

(As+b)(s+1)(s+2)+ c (s^2)(s+2) + d s^2 (s+1) =(2+s^2)

OpenStudy (anonymous):

thanks! I think I got it.

OpenStudy (across):

Are you working with Laplace transforms? Break the quotient as follows: \[\frac{2+s^2}{s^2(s+1)(s+2)}=\frac{As+D}{s^2}+\frac{B}{s+1}+\frac{C}{s+2}.\]Now it's just a matter of solving for A, B and C. \[(As+D)(s+1)(s+2)+Bs^2(s+2)+Cs^2(s+1)=2+s^2.\]Setting s = -1, \[B=3.\]Setting s = -2,\[-4C=6\implies C=-\frac{3}{2}.\]Setting s = 0,\[2D=2\implies D=1.\]Finally, \[A=-\frac{3}{2}.\]Therefore,\[\frac{1}{s^2}+\frac{3}{s+1}-\frac{3}{2(s+2)}-\frac{3}{2s}.\]

OpenStudy (across):

Too slow. xd

OpenStudy (anonymous):

yes, I am :) fun stuff lol.